Lisa bought on calculator and a textbook in math. Her bill including the 7.25% sales tax came out to be $100.58. The textbook was five dollars less than twice the price of the calculator.
How much was each item before tax?

Answer by rfer(16322)

2x-5+x=100.58*0.9275
3x-5=98.07
3x=103.07
x=34.36
2x-5=63.71